OpenStudy (anonymous):

Two rubber bands pulling on an object cause it to accelerate at 1.2 m/s^2. a. What will be the object's acceleration if it is pulled by four rubber bands? b. What will be the acceleration of two of these objects glued together if they are pulled be two rubber bands?

OpenStudy (plainntall):

Assume the rubber bands have no mass and are identical so they supply the same force to the object. Newton's Second Law of motion says F=ma where F is the net force acting on the object. solve for a

OpenStudy (anonymous):

a=F/m

OpenStudy (anonymous):

now what?

OpenStudy (anonymous):

hello

OpenStudy (plainntall):

Well what happens when you double the amount of force and keep all other values the same?

OpenStudy (plainntall):

2F/m=2a=2(1.2m/s^2) ; you double the acceleration correct?

OpenStudy (anonymous):

so this is proportionality question.

OpenStudy (plainntall):

yeah which is why question b askes what happens if you double the mass

OpenStudy (plainntall):

a) what happens to acceleration when you double the force b) what happens to acceleration when you double the mass

OpenStudy (anonymous):

b then gonna be F/2m=a

OpenStudy (plainntall):

so doubling the mass halves the acceleration right?

OpenStudy (plainntall):

so a) 2(1.2) b) 1.2/2

OpenStudy (anonymous):

symbolically dose not make sense to me. For both a and b.

OpenStudy (anonymous):

I think I go it now thank very much.

OpenStudy (plainntall):

welcome